### v3

# The T2K Collaboration Results from the T2K experiment on neutrino mixing including a new far detector -like sample K. Abe

††thanks: also at Université Paris-Saclay††thanks: also at Departament de Fisica de la Universitat Autonoma de Barcelona, Barcelona, Spain††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at Moscow Institute of Physics and Technology (MIPT), Moscow region, Russia and National Research Nuclear University “MEPhI”, Moscow, Russia††thanks: also at IPSA-DRII, France††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at JINR, Dubna, Russia††thanks: also at the Graduate University of Science and Technology, Vietnam Academy of Science and Technology††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at Nambu Yoichiro Institute of Theoretical and Experimental Physics (NITEP)††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at Nambu Yoichiro Institute of Theoretical and Experimental Physics (NITEP)††thanks: also at BMCC/CUNY, Science Department, New York, New York, U.S.A.††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan
The T2K Collaboration
Results from the T2K experiment on neutrino mixing including a new far detector -like sample
K. Abe
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
S. Abe
University of Tokyo, Department of Physics, Tokyo, Japan
R. Akutsu
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
H. Alarakia-Charles
Lancaster University, Physics Department, Lancaster, United Kingdom
Y.I. Alj Hakim
University of Sheffield, School of Mathematical and Physical Sciences, Sheffield, United Kingdom
S. Alonso Monsalve
ETH Zurich, Institute for Particle Physics and Astrophysics, Zurich, Switzerland
L. Anthony
Imperial College London, Department of Physics, London, United Kingdom
S. Aoki
Kobe University, Kobe, Japan
K.A. Apte
Imperial College London, Department of Physics, London, United Kingdom
T. Arai
University of Tokyo, Department of Physics, Tokyo, Japan
T. Arihara
Tokyo Metropolitan University, Department of Physics, Tokyo, Japan
S. Arimoto
Kyoto University, Department of Physics, Kyoto, Japan
Y. Ashida
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
E.T. Atkin
Imperial College London, Department of Physics, London, United Kingdom
N. Babu
Louisiana State University, Department of Physics and Astronomy, Baton Rouge, Louisiana, U.S.A.
V. Baranov
Joint Institute for Nuclear Research, Dubna, Moscow Region, Russia
G.J. Barker
University of Warwick, Department of Physics, Coventry, United Kingdom
G. Barr
Oxford University, Department of Physics, Oxford, United Kingdom
D. Barrow
Oxford University, Department of Physics, Oxford, United Kingdom
P. Bates
University of Liverpool, Department of Physics, Liverpool, United Kingdom
L. Bathe-Peters
Oxford University, Department of Physics, Oxford, United Kingdom
M. Batkiewicz-Kwasniak
H. Niewodniczanski Institute of Nuclear Physics PAN, Cracow, Poland
N. Baudis
Oxford University, Department of Physics, Oxford, United Kingdom
V. Berardi
INFN Sezione di Bari and Università e Politecnico di Bari, Dipartimento Interuniversitario di Fisica, Bari, Italy
L. Berns
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
S. Bhattacharjee
Louisiana State University, Department of Physics and Astronomy, Baton Rouge, Louisiana, U.S.A.
A. Blanchet
CERN European Organization for Nuclear Research, CH-1211 Genéve 23, Switzerland
A. Blondel
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
P.M.M. Boistier
IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
S. Bolognesi
IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
S. Bordoni
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
S.B. Boyd
University of Warwick, Department of Physics, Coventry, United Kingdom
C. Bronner
Yokohama National University, Department of Physics, Yokohama, Japan
A. Bubak
University of Silesia, Institute of Physics, Katowice, Poland
M. Buizza Avanzini
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
J.A. Caballero
Departamento de Física Atómica, Molecular y Nuclear, Universidad de Sevilla, 41080 Sevilla, Spain
F. Cadoux
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
N.F. Calabria
INFN Sezione di Bari and Università e Politecnico di Bari, Dipartimento Interuniversitario di Fisica, Bari, Italy
S. Cao
Institute For Interdisciplinary Research in Science and Education (IFIRSE), ICISE, Quy Nhon, Vietnam
S. Cap
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
D. Carabadjac
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
S.L. Cartwright
University of Sheffield, School of Mathematical and Physical Sciences, Sheffield, United Kingdom
M.P. Casado
Institut de Fisica d’Altes Energies (IFAE) - The Barcelona Institute of Science and Technology, Campus UAB, Bellaterra (Barcelona) Spain
M.G. Catanesi
INFN Sezione di Bari and Università e Politecnico di Bari, Dipartimento Interuniversitario di Fisica, Bari, Italy
J. Chakrani
Lawrence Berkeley National Laboratory, Berkeley, California, U.S.A.
A. Chalumeau
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
D. Cherdack
University of Houston, Department of Physics, Houston, Texas, U.S.A.
A. Chvirova
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
J. Coleman
University of Liverpool, Department of Physics, Liverpool, United Kingdom
G. Collazuol
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
F. Cormier
TRIUMF, Vancouver, British Columbia, Canada
A.A.L. Craplet
Imperial College London, Department of Physics, London, United Kingdom
A. Cudd
University of Colorado at Boulder, Department of Physics, Boulder, Colorado, U.S.A.
D. D’Ago
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
C. Dalmazzone
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
T. Daret
IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
P. Dasgupta
Eötvös Loránd University, Department of Atomic Physics, Budapest, Hungary
C. Davis
University of Pennsylvania, Department of Physics and Astronomy, Philadelphia, Pennsylvania, U.S.A.
Yu.I. Davydov
Joint Institute for Nuclear Research, Dubna, Moscow Region, Russia
P. de Perio
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
G. De Rosa
INFN Sezione di Napoli and Università di Napoli, Dipartimento di Fisica, Napoli, Italy
T. Dealtry
Lancaster University, Physics Department, Lancaster, United Kingdom
C. Densham
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
A. Dergacheva
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
R. Dharmapal Banerjee
Wroclaw University, Faculty of Physics and Astronomy, Wroclaw, Poland
F. Di Lodovico
King’s College London, Department of Physics, Strand, London WC2R 2LS, United Kingdom
G. Diaz Lopez
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
S. Dolan
CERN European Organization for Nuclear Research, CH-1211 Genéve 23, Switzerland
D. Douqa
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
T.A. Doyle
Oxford University, Department of Physics, Oxford, United Kingdom
O. Drapier
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
K.E. Duffy
Oxford University, Department of Physics, Oxford, United Kingdom
J. Dumarchez
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
P. Dunne
Imperial College London, Department of Physics, London, United Kingdom
K. Dygnarowicz
Warsaw University of Technology, Institute of Radioelectronics and Multimedia Technology, Warsaw, Poland
A. Eguchi
University of Tokyo, Department of Physics, Tokyo, Japan
J. Elias
University of Rochester, Department of Physics and Astronomy, Rochester, New York, U.S.A.
S. Emery-Schrenk
IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
G. Erofeev
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
A. Ershova
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
G. Eurin
IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
D. Fedorova
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
S. Fedotov
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
M. Feltre
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
L. Feng
Kyoto University, Department of Physics, Kyoto, Japan
D. Ferlewicz
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
A.J. Finch
Lancaster University, Physics Department, Lancaster, United Kingdom
M.D. Fitton
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
C. Forza
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
M. Friend
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
Y. Fujii
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
Y. Fukuda
Miyagi University of Education, Department of Physics, Sendai, Japan
Y. Furui
Tokyo Metropolitan University, Department of Physics, Tokyo, Japan
J. García-Marcos
Ghent University, Department of Physics and Astronomy, Proeftuinstraat 86, B-9000 Gent, Belgium
A.C. Germer
University of Pennsylvania, Department of Physics and Astronomy, Philadelphia, Pennsylvania, U.S.A.
L. Giannessi
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
C. Giganti
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
M. Girgus
University of Warsaw, Faculty of Physics, Warsaw, Poland
V. Glagolev
Joint Institute for Nuclear Research, Dubna, Moscow Region, Russia
M. Gonin
ILANCE, CNRS – University of Tokyo International Research Laboratory, Kashiwa, Chiba 277-8582, Japan
R. González Jiménez
Departamento de Física Atómica, Molecular y Nuclear, Universidad de Sevilla, 41080 Sevilla, Spain
J. González Rosa
Departamento de Física Atómica, Molecular y Nuclear, Universidad de Sevilla, 41080 Sevilla, Spain
E.A.G. Goodman
University of Glasgow, School of Physics and Astronomy, Glasgow, United Kingdom
K. Gorshanov
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
P. Govindaraj
University of Warsaw, Faculty of Physics, Warsaw, Poland
M. Grassi
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
M. Guigue
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
F.Y. Guo
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
D.R. Hadley
University of Warwick, Department of Physics, Coventry, United Kingdom
S. Han
Kyoto University, Department of Physics, Kyoto, Japan
University of Tokyo, Institute for Cosmic Ray Research, Research Center for Cosmic Neutrinos, Kashiwa, Japan
D.A. Harris
York University, Department of Physics and Astronomy, Toronto, Ontario, Canada
R.J. Harris
Lancaster University, Physics Department, Lancaster, United Kingdom
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
T. Hasegawa
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
C.M. Hasnip
CERN European Organization for Nuclear Research, CH-1211 Genéve 23, Switzerland
S. Hassani
IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
N.C. Hastings
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
Y. Hayato
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
I. Heitkamp
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
D. Henaff
IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
Y. Hino
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
J. Holeczek
University of Silesia, Institute of Physics, Katowice, Poland
A. Holin
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
T. Holvey
Oxford University, Department of Physics, Oxford, United Kingdom
N.T. Hong Van
International Centre of Physics, Institute of Physics (IOP), Vietnam Academy of Science and Technology (VAST), 10 Dao Tan, Ba Dinh, Hanoi, Vietnam
T. Honjo
Osaka Metropolitan University, Department of Physics, Osaka, Japan
M.C.F. Hooft
Ghent University, Department of Physics and Astronomy, Proeftuinstraat 86, B-9000 Gent, Belgium
K. Hosokawa
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
J. Hu
Kyoto University, Department of Physics, Kyoto, Japan
A.K. Ichikawa
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
K. Ieki
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
M. Ikeda
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
T.H. Ishida
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
T. Ishida
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
M. Ishitsuka
Tokyo University of Science, Faculty of Science and Technology, Department of Physics, Noda, Chiba, Japan
H. Ito
Kobe University, Kobe, Japan
S. Ito
Yokohama National University, Department of Physics, Yokohama, Japan
A. Izmaylov
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
N. Jachowicz
Ghent University, Department of Physics and Astronomy, Proeftuinstraat 86, B-9000 Gent, Belgium
S.J. Jenkins
University of Liverpool, Department of Physics, Liverpool, United Kingdom
C. Jesús-Valls
CERN European Organization for Nuclear Research, CH-1211 Genéve 23, Switzerland
M. Jia
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
J.J. Jiang
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
J.Y. Ji
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
T.P. Jones
Lancaster University, Physics Department, Lancaster, United Kingdom
P. Jonsson
Imperial College London, Department of Physics, London, United Kingdom
S. Joshi
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
C.K. Jung
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
M. Kabirnezhad
Imperial College London, Department of Physics, London, United Kingdom
A.C. Kaboth
Royal Holloway University of London, Department of Physics, Egham, Surrey, United Kingdom
H. Kakuno
Tokyo Metropolitan University, Department of Physics, Tokyo, Japan
J. Kameda
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
S. Karpova
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
V.S. Kasturi
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
Y. Kataoka
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
T. Katori
King’s College London, Department of Physics, Strand, London WC2R 2LS, United Kingdom
A. Kawabata
Keio University, Department of Physics, Kanagawa, Japan
Y. Kawamura
Osaka Metropolitan University, Department of Physics, Osaka, Japan
M. Kawaue
Kyoto University, Department of Physics, Kyoto, Japan
E. Kearns
Boston University, Department of Physics, Boston, Massachusetts, U.S.A.
M. Khabibullin
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
A. Khotjantsev
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
T. Kikawa
Kyoto University, Department of Physics, Kyoto, Japan
S. King
King’s College London, Department of Physics, Strand, London WC2R 2LS, United Kingdom
V. Kiseeva
Joint Institute for Nuclear Research, Dubna, Moscow Region, Russia
J. Kisiel
University of Silesia, Institute of Physics, Katowice, Poland
A. Klustová
Imperial College London, Department of Physics, London, United Kingdom
L. Kneale
University of Sheffield, School of Mathematical and Physical Sciences, Sheffield, United Kingdom
H. Kobayashi
University of Tokyo, Department of Physics, Tokyo, Japan
Sota.R Kobayashi
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
L. Koch
Institut für Physik, Johannes Gutenberg-Universität Mainz, Staudingerweg 7, 55128 Mainz, Germany
S. Kodama
University of Tokyo, Department of Physics, Tokyo, Japan
M. Kolupanova
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
A. Konaka
TRIUMF, Vancouver, British Columbia, Canada
L.L. Kormos
Lancaster University, Physics Department, Lancaster, United Kingdom
Y. Koshio
Okayama University, Department of Physics, Okayama, Japan
K. Kowalik
National Centre for Nuclear Research, Warsaw, Poland
Y. Kudenko
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
Y. Kudo
Yokohama National University, Department of Physics, Yokohama, Japan
A. Kumar Jha
Ghent University, Department of Physics and Astronomy, Proeftuinstraat 86, B-9000 Gent, Belgium
R. Kurjata
Warsaw University of Technology, Institute of Radioelectronics and Multimedia Technology, Warsaw, Poland
V. Kurochka
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
T. Kutter
Louisiana State University, Department of Physics and Astronomy, Baton Rouge, Louisiana, U.S.A.
L. Labarga
University Autonoma Madrid, Department of Theoretical Physics, 28049 Madrid, Spain
M. Lachat
University of Rochester, Department of Physics and Astronomy, Rochester, New York, U.S.A.
K. Lachner
ETH Zurich, Institute for Particle Physics and Astrophysics, Zurich, Switzerland
J. Lagoda
National Centre for Nuclear Research, Warsaw, Poland
S.M. Lakshmi
University of Silesia, Institute of Physics, Katowice, Poland
M. Lamers James
University of Warwick, Department of Physics, Coventry, United Kingdom
A. Langella
INFN Sezione di Napoli and Università di Napoli, Dipartimento di Fisica, Napoli, Italy
D.H. Langridge
Royal Holloway University of London, Department of Physics, Egham, Surrey, United Kingdom
J.-F. Laporte
IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
D. Last
University of Rochester, Department of Physics and Astronomy, Rochester, New York, U.S.A.
N. Latham
King’s College London, Department of Physics, Strand, London WC2R 2LS, United Kingdom
M. Laveder
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
L. Lavitola
INFN Sezione di Napoli and Università di Napoli, Dipartimento di Fisica, Napoli, Italy
M. Lawe
Lancaster University, Physics Department, Lancaster, United Kingdom
D. Leon Silverio
South Dakota School of Mines and Technology, 501 East Saint Joseph Street, Rapid City, SD 57701, United States
S. Levorato
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
S.V. Lewis
King’s College London, Department of Physics, Strand, London WC2R 2LS, United Kingdom
B. Li
ETH Zurich, Institute for Particle Physics and Astrophysics, Zurich, Switzerland
C. Lin
Imperial College London, Department of Physics, London, United Kingdom
R.P. Litchfield
University of Glasgow, School of Physics and Astronomy, Glasgow, United Kingdom
S.L. Liu
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
W. Li
Oxford University, Department of Physics, Oxford, United Kingdom
A. Longhin
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
A. Lopez Moreno
King’s College London, Department of Physics, Strand, London WC2R 2LS, United Kingdom
L. Ludovici
INFN Sezione di Roma and Università di Roma “La Sapienza”, Roma, Italy
X. Lu
University of Warwick, Department of Physics, Coventry, United Kingdom
T. Lux
Institut de Fisica d’Altes Energies (IFAE) - The Barcelona Institute of Science and Technology, Campus UAB, Bellaterra (Barcelona) Spain
L.N. Machado
University of Glasgow, School of Physics and Astronomy, Glasgow, United Kingdom
L. Magaletti
INFN Sezione di Bari and Università e Politecnico di Bari, Dipartimento Interuniversitario di Fisica, Bari, Italy
K. Mahn
Michigan State University, Department of Physics and Astronomy, East Lansing, Michigan, U.S.A.
K.K. Mahtani
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
M. Mandal
National Centre for Nuclear Research, Warsaw, Poland
S. Manly
University of Rochester, Department of Physics and Astronomy, Rochester, New York, U.S.A.
A.D. Marino
University of Colorado at Boulder, Department of Physics, Boulder, Colorado, U.S.A.
D.G.R. Martin
Imperial College London, Department of Physics, London, United Kingdom
D.A. Martinez Caicedo
South Dakota School of Mines and Technology, 501 East Saint Joseph Street, Rapid City, SD 57701, United States
L. Martinez
Institut de Fisica d’Altes Energies (IFAE) - The Barcelona Institute of Science and Technology, Campus UAB, Bellaterra (Barcelona) Spain
M. Martini
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
T. Matsubara
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
R. Matsumoto
Institute of Science Tokyo, Department of Physics, Tokyo
V. Matveev
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
C. Mauger
University of Pennsylvania, Department of Physics and Astronomy, Philadelphia, Pennsylvania, U.S.A.
K. Mavrokoridis
University of Liverpool, Department of Physics, Liverpool, United Kingdom
N. McCauley
University of Liverpool, Department of Physics, Liverpool, United Kingdom
K.S. McFarland
University of Rochester, Department of Physics and Astronomy, Rochester, New York, U.S.A.
C. McGrew
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
J. McKean
Imperial College London, Department of Physics, London, United Kingdom
A. Mefodiev
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
G.D. Megias
Departamento de Física Atómica, Molecular y Nuclear, Universidad de Sevilla, 41080 Sevilla, Spain
L. Mellet
Michigan State University, Department of Physics and Astronomy, East Lansing, Michigan, U.S.A.
C. Metelko
University of Liverpool, Department of Physics, Liverpool, United Kingdom
M. Mezzetto
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
S. Miki
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
V. Mikola
University of Glasgow, School of Physics and Astronomy, Glasgow, United Kingdom
E.W. Miller
Imperial College London, Department of Physics, London, United Kingdom
A. Minamino
Yokohama National University, Department of Physics, Yokohama, Japan
O. Mineev
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
S. Mine
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
University of California, Irvine, Department of Physics and Astronomy, Irvine, California, U.S.A.
J. Mirabito
Boston University, Department of Physics, Boston, Massachusetts, U.S.A.
M. Miura
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
S. Moriyama
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
S. Moriyama
Yokohama National University, Department of Physics, Yokohama, Japan
P. Morrison
University of Glasgow, School of Physics and Astronomy, Glasgow, United Kingdom
Th.A. Mueller
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
D. Munford
University of Houston, Department of Physics, Houston, Texas, U.S.A.
A. Muñoz
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
ILANCE, CNRS – University of Tokyo International Research Laboratory, Kashiwa, Chiba 277-8582, Japan
L. Munteanu
CERN European Organization for Nuclear Research, CH-1211 Genéve 23, Switzerland
Y. Nagai
Eötvös Loránd University, Department of Atomic Physics, Budapest, Hungary
T. Nakadaira
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
K. Nakagiri
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
M. Nakahata
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
Y. Nakajima
University of Tokyo, Department of Physics, Tokyo, Japan
K.D. Nakamura
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
A. Nakano
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
Y. Nakano
University of Toyama, Department of Physics, Toyama, Japan
S. Nakayama
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
T. Nakaya
Kyoto University, Department of Physics, Kyoto, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
K. Nakayoshi
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
C.E.R. Naseby
Imperial College London, Department of Physics, London, United Kingdom
D.T. Nguyen
VNU University of Science, Vietnam National University, Hanoi, Vietnam
V.Q. Nguyen
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
K. Niewczas
Ghent University, Department of Physics and Astronomy, Proeftuinstraat 86, B-9000 Gent, Belgium
S. Nishimori
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
Y. Nishimura
Keio University, Department of Physics, Kanagawa, Japan
Y. Noguchi
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
T. Nosek
National Centre for Nuclear Research, Warsaw, Poland
F. Nova
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
J.C. Nugent
Imperial College London, Department of Physics, London, United Kingdom
H.M. O’Keeffe
Lancaster University, Physics Department, Lancaster, United Kingdom
L. O’Sullivan
Institut für Physik, Johannes Gutenberg-Universität Mainz, Staudingerweg 7, 55128 Mainz, Germany
R. Okazaki
Keio University, Department of Physics, Kanagawa, Japan
W. Okinaga
University of Tokyo, Department of Physics, Tokyo, Japan
K. Okumura
University of Tokyo, Institute for Cosmic Ray Research, Research Center for Cosmic Neutrinos, Kashiwa, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
T. Okusawa
Osaka Metropolitan University, Department of Physics, Osaka, Japan
N. Onda
Kyoto University, Department of Physics, Kyoto, Japan
N. Ospina
INFN Sezione di Bari and Università e Politecnico di Bari, Dipartimento Interuniversitario di Fisica, Bari, Italy
L. Osu
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
N. Otani
Kyoto University, Department of Physics, Kyoto, Japan
Y. Oyama
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
V. Paolone
University of Pittsburgh, Department of Physics and Astronomy, Pittsburgh, Pennsylvania, U.S.A.
J. Pasternak
Imperial College London, Department of Physics, London, United Kingdom
D. Payne
University of Liverpool, Department of Physics, Liverpool, United Kingdom
T.P.D. Peacock
University of Sheffield, School of Mathematical and Physical Sciences, Sheffield, United Kingdom
M. Pfaff
Imperial College London, Department of Physics, London, United Kingdom
L. Pickering
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
B. Popov
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
A.J. Portocarrero Yrey
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
M. Posiadala-Zezula
University of Warsaw, Faculty of Physics, Warsaw, Poland
Y.S. Prabhu
University of Warsaw, Faculty of Physics, Warsaw, Poland
H. Prasad
Wroclaw University, Faculty of Physics and Astronomy, Wroclaw, Poland
F. Pupilli
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
B. Quilain
ILANCE, CNRS – University of Tokyo International Research Laboratory, Kashiwa, Chiba 277-8582, Japan
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
P.T. Quyen
Institute For Interdisciplinary Research in Science and Education (IFIRSE), ICISE, Quy Nhon, Vietnam
E. Radicioni
INFN Sezione di Bari and Università e Politecnico di Bari, Dipartimento Interuniversitario di Fisica, Bari, Italy
B. Radics
York University, Department of Physics and Astronomy, Toronto, Ontario, Canada
M.A. Ramirez Delgado
University of Pennsylvania, Department of Physics and Astronomy, Philadelphia, Pennsylvania, U.S.A.
R. Ramsden
King’s College London, Department of Physics, Strand, London WC2R 2LS, United Kingdom
P.N. Ratoff
Lancaster University, Physics Department, Lancaster, United Kingdom
M. Reh
University of Colorado at Boulder, Department of Physics, Boulder, Colorado, U.S.A.
G. Reina
Institut für Physik, Johannes Gutenberg-Universität Mainz, Staudingerweg 7, 55128 Mainz, Germany
C. Riccio
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
D.W. Riley
University of Glasgow, School of Physics and Astronomy, Glasgow, United Kingdom
E. Rondio
National Centre for Nuclear Research, Warsaw, Poland
S. Roth
RWTH Aachen University, III. Physikalisches Institut, Aachen, Germany
N. Roy
York University, Department of Physics and Astronomy, Toronto, Ontario, Canada
A. Rubbia
ETH Zurich, Institute for Particle Physics and Astrophysics, Zurich, Switzerland
L. Russo
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
A. Rychter
Warsaw University of Technology, Institute of Radioelectronics and Multimedia Technology, Warsaw, Poland
W. Saenz
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
K. Sakashita
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
S. Samani
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
F. Sánchez
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
E.M. Sandford
University of Liverpool, Department of Physics, Liverpool, United Kingdom
Y. Sato
Tokyo University of Science, Faculty of Science and Technology, Department of Physics, Noda, Chiba, Japan
T. Schefke
Louisiana State University, Department of Physics and Astronomy, Baton Rouge, Louisiana, U.S.A.
C.M. Schloesser
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
K. Scholberg
Duke University, Department of Physics, Durham, North Carolina, U.S.A.
M. Scott
Imperial College London, Department of Physics, London, United Kingdom
Y. Seiya
Osaka Metropolitan University, Department of Physics, Osaka, Japan
T. Sekiguchi
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
H. Sekiya
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
T. Sekiya
Tokyo Metropolitan University, Department of Physics, Tokyo, Japan
D. Seppala
Michigan State University, Department of Physics and Astronomy, East Lansing, Michigan, U.S.A.
D. Sgalaberna
ETH Zurich, Institute for Particle Physics and Astrophysics, Zurich, Switzerland
A. Shaikhiev
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
M. Shiozawa
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
Y. Shiraishi
Okayama University, Department of Physics, Okayama, Japan
A. Shvartsman
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
N. Skrobova
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
K. Skwarczynski
Royal Holloway University of London, Department of Physics, Egham, Surrey, United Kingdom
D. Smyczek
RWTH Aachen University, III. Physikalisches Institut, Aachen, Germany
M. Smy
University of California, Irvine, Department of Physics and Astronomy, Irvine, California, U.S.A.
J.T. Sobczyk
Wroclaw University, Faculty of Physics and Astronomy, Wroclaw, Poland
H. Sobel
University of California, Irvine, Department of Physics and Astronomy, Irvine, California, U.S.A.
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
F.J.P. Soler
University of Glasgow, School of Physics and Astronomy, Glasgow, United Kingdom
A.J. Speers
Lancaster University, Physics Department, Lancaster, United Kingdom
R. Spina
INFN Sezione di Bari and Università e Politecnico di Bari, Dipartimento Interuniversitario di Fisica, Bari, Italy
A. Srivastava
Institut für Physik, Johannes Gutenberg-Universität Mainz, Staudingerweg 7, 55128 Mainz, Germany
P. Stowell
University of Sheffield, School of Mathematical and Physical Sciences, Sheffield, United Kingdom
Y. Stroke
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
I.A. Suslov
Joint Institute for Nuclear Research, Dubna, Moscow Region, Russia
A. Suzuki
Kobe University, Kobe, Japan
S.Y. Suzuki
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
M. Tada
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
S. Tairafune
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
A. Takeda
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Y. Takeuchi
Kobe University, Kobe, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
K. Takeya
Okayama University, Department of Physics, Okayama, Japan
H.K. Tanaka
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
H. Tanigawa
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
A. Teklu
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
V.V. Tereshchenko
Joint Institute for Nuclear Research, Dubna, Moscow Region, Russia
N. Thamm
RWTH Aachen University, III. Physikalisches Institut, Aachen, Germany
C. Touramanis
University of Liverpool, Department of Physics, Liverpool, United Kingdom
N. Tran
Kyoto University, Department of Physics, Kyoto, Japan
T. Tsukamoto
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
M. Tzanov
Louisiana State University, Department of Physics and Astronomy, Baton Rouge, Louisiana, U.S.A.
Y. Uchida
Imperial College London, Department of Physics, London, United Kingdom
M. Vagins
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
University of California, Irvine, Department of Physics and Astronomy, Irvine, California, U.S.A.
M. Varghese
Institut de Fisica d’Altes Energies (IFAE) - The Barcelona Institute of Science and Technology, Campus UAB, Bellaterra (Barcelona) Spain
I. Vasilyev
Joint Institute for Nuclear Research, Dubna, Moscow Region, Russia
G. Vasseur
IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
E. Villa
CERN European Organization for Nuclear Research, CH-1211 Genéve 23, Switzerland
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
U. Virginet
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
T. Vladisavljevic
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
T. Wachala
H. Niewodniczanski Institute of Nuclear Physics PAN, Cracow, Poland
S.-i. Wada
Kobe University, Kobe, Japan
D. Wakabayashi
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
H.T. Wallace
University of Sheffield, School of Mathematical and Physical Sciences, Sheffield, United Kingdom
J.G. Walsh
Michigan State University, Department of Physics and Astronomy, East Lansing, Michigan, U.S.A.
L. Wan
Boston University, Department of Physics, Boston, Massachusetts, U.S.A.
D. Wark
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
Oxford University, Department of Physics, Oxford, United Kingdom
M.O. Wascko
Oxford University, Department of Physics, Oxford, United Kingdom
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
A. Weber
Institut für Physik, Johannes Gutenberg-Universität Mainz, Staudingerweg 7, 55128 Mainz, Germany
R. Wendell
Kyoto University, Department of Physics, Kyoto, Japan
M.J. Wilking
School of Physics and Astronomy, University of Minnesota, Minneapolis, Minnesota, U.S.A.
C. Wilkinson
Lawrence Berkeley National Laboratory, Berkeley, California, U.S.A.
J.R. Wilson
King’s College London, Department of Physics, Strand, London WC2R 2LS, United Kingdom
K. Wood
Lawrence Berkeley National Laboratory, Berkeley, California, U.S.A.
C. Wret
Imperial College London, Department of Physics, London, United Kingdom
J. Xia
SLAC National Accelerator Laboratory, Stanford University, Menlo Park, California, U.S.A.
K. Yamamoto
Osaka Metropolitan University, Department of Physics, Osaka, Japan
T. Yamamoto
Osaka Metropolitan University, Department of Physics, Osaka, Japan
C. Yanagisawa
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
Y. Yang
Oxford University, Department of Physics, Oxford, United Kingdom
T. Yano
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
K. Yasutome
Kyoto University, Department of Physics, Kyoto, Japan
N. Yershov
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
U. Yevarouskaya
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
M. Yokoyama
University of Tokyo, Department of Physics, Tokyo, Japan
Y. Yoshimoto
University of Tokyo, Department of Physics, Tokyo, Japan
N. Yoshimura
Kyoto University, Department of Physics, Kyoto, Japan
R. Zaki
York University, Department of Physics and Astronomy, Toronto, Ontario, Canada
A. Zalewska
H. Niewodniczanski Institute of Nuclear Physics PAN, Cracow, Poland
J. Zalipska
National Centre for Nuclear Research, Warsaw, Poland
G. Zarnecki
H. Niewodniczanski Institute of Nuclear Physics PAN, Cracow, Poland
J. Zhang
TRIUMF, Vancouver, British Columbia, Canada
University of British Columbia, Department of Physics and Astronomy, Vancouver, British Columbia, Canada
X.Y. Zhao
ETH Zurich, Institute for Particle Physics and Astrophysics, Zurich, Switzerland
H. Zheng
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
H. Zhong
Kobe University, Kobe, Japan
T. Zhu
Imperial College London, Department of Physics, London, United Kingdom
M. Ziembicki
Warsaw University of Technology, Institute of Radioelectronics and Multimedia Technology, Warsaw, Poland
E.D. Zimmerman
University of Colorado at Boulder, Department of Physics, Boulder, Colorado, U.S.A.
M. Zito
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
S. Zsoldos
King’s College London, Department of Physics, Strand, London WC2R 2LS, United Kingdom
(December 30, 2025)
Abstract
We have made improved measurements of three-flavor neutrino mixing with 19.7(16.3) protons on target in (anti-)neutrino-enhanced beam modes. A new sample of muon-neutrino events with tagged pions has been added at the far detector, as well as new proton and photon-tagged samples at the near detector. Significant improvements have been made to the flux and neutrino interaction modeling. T2K data continue to prefer the normal mass ordering and upper octant of with a near-maximal value of the charge-parity violating phase with best-fit values in the normal ordering of , and eV2.
Introduction— The T2K experiment [1] measures three-flavor neutrino mixing parameters by observing the disappearance of () and the appearance of () over a distance of 295 km in a narrow-band, predominantly or beam, which peaks at an energy of 0.6 GeV. Data corresponding to 3.61021 protons on target (POT) were analyzed with major improvements to the neutrino flux and interaction modeling compared to previous studies [2]. New photon- and proton-tagged near detector (ND) samples, as well as a new, charged current neutrino-enhanced muon-neutrino sample with tagged pions (CC-like) in the far detector, are included in the analysis for the first time. This letter reports T2K’s latest measurements of neutrino mixing parameters after these improvements, which mark a significant step forward in the robustness of our analysis, particularly due to improvements in modeling and constraining neutrino interaction uncertainties, and supersede previous results produced with the same dataset.
Comparable measurements have also been reported by NOvA [3], and future experiments such as Hyper-Kamiokande [4] and DUNE [5] aim to further improve the precision and discovery potential in neutrino oscillation physics.
The T2K experiment—
Protons of 30 GeV energy, accelerated by the J-PARC main ring, collide and interact with a graphite target, producing pions and kaons that are focused by a system of three magnetic horns and decay inside a 96 m long tunnel. Depending on the direction of the current flowing in the horns, either positively or negatively charged hadrons are focused, producing a neutrino- (-mode) or antineutrino-enhanced (-mode) beam.
Two near detectors are located 280 m away from the graphite target — one on-axis (INGRID) [6], and the other 2.5∘ off-axis (ND280) [1] with respect to the beam direction. They sample the unoscillated beam by monitoring its direction, intensity and flavor content as well as constraining uncertainties in the neutrino interaction model. The far detector, Super-Kamiokande (SK) [7], is a 50 kton water-Cherenkov detector located beneath a 1 km rock overburden within the Kamioka mine in Japan. It measures the oscillated neutrino flux 295 km away from its production point at off-axis.
The analysis reported in this letter uses all data collected by T2K from January 2010 to February 2020, corresponding to an SK exposure of POT in -mode and POT in -mode. The same number of POT was used in a previous analysis [2].
Neutrino flux prediction— The neutrino flux Monte Carlo (MC) simulation uses the FLUKA 2011.2 [8, 9] interaction model for proton interactions inside the target and GEANT3 [10] for interactions outside the target. Hadron production in the target is tuned to external datasets [11, 12, 13]. The flux depends on the beamline conditions and the measured proton beam profile. The prediction accounts for the beam conditions during each operating period.
The resulting flux model [14] is used to estimate the unoscillated -mode and -mode fluxes at all detectors, for all contributing neutrino flavors, as well as their correlated uncertainties. INGRID is used to monitor the beam direction and validate the neutrino flux simulation. The uncertainty on the measured beam direction is included in the flux uncertainties used in the oscillation parameter measurement.
This analysis includes new constraints from the NA61/SHINE replica target datasets [13], including high-statistics measurements of production as well as, for the first time, and proton production measurements. Among other improvements these constraints reduce the flux uncertainty to below 4% for energies up to 7 GeV.
Neutrino interaction modeling— The nominal model prediction was generated with the neutrino event generator NEUT 5.4.0 [15]. In this analysis, important improvements to the estimated neutrino interaction uncertainties were made, increasing the number of free parameters by 26 compared with the previous analysis [2], for a total of 75.
Pionless:
New degrees of freedom were included to cover theoretical uncertainties in nuclear effects for Charged-Current (CC) Quasi-Elastic (CCQE), multi-nucleon (2p2h) and hard-scattering processes.
Ad hoc uncertainties in the low four-momentum transfer () response for CCQE interactions were replaced with physics motivated parameters that vary the impact of Pauli blocking and the effect of nuclear transparency on the inclusive cross section. An effective parameterization of the latter was derived by comparison to the NuWro [16, 17] implementation of the optical potential [18]. New uncertainties on the nuclear ground state for CCQE events, which parameterize the shell structure of the missing momentum () and missing energy () response in the initial state Spectral Function (SF), were included [19, 20]. The Benhar SF [21] incorporates multi-nucleon knockout in the MeV, MeV region. These events account for 5% of all CCQE interactions; a 200% normalization uncertainty was applied to cover differences between NEUT and NuWro predictions. Following Ref. [22], a three-momentum-transfer-dependent freedom on the Nucleon Removal Energy (NRE) was added [23]. Additional freedom for 2p2h interactions was included by separating some of the existing uncertainties by the struck nucleon pair (proton–neutron vs. proton–proton/neutron–neutron).
Single pion production (SPP):
Uncertainties on the free parameters in the NEUT implementation [2] of the Rein–Sehgal SPP model [24] were motivated by fitting [19] to hydrogen- and deuterium-target data [25, 26] from ANL [27, 28] and BNL [29, 30, 31], and checking coverage against hydrocarbon-target data [32, 33, 34, 35, 36, 37, 38]. Variations of these parameters exhibit a minimal change to both the shape of the final-state pion kinematics, which have previously been observed to be poorly predicted at both ND280 and SK [2], and to the relative rate of charged and neutral pion production. To cover measurements of pion kinematic spectra in T2K [39] and MINERA [40], new parameters were developed to vary the resonance decay kinematics and the total neutral pion production rate [19]. Previously, NRE effects were only included in NEUT for pionless hard-scatter processes. Here, an approximation of this effect, and an associated uncertainty, was motivated by comparing NEUT predictions to those from NuWro for MeV.
Multiple pions:
New uncertainties were developed for multiple pion production in the low hadronic mass region, . Ref. [41] motivated separate uncertainties on the axial and vector parts of the nucleon form factors, replacing an effective 100%, -dependent, uncertainty on the previous Bodek-Yang correction [42]. Additionally, a new uncertainty covering the difference between the NEUT and AGKY models [43] was included on the shape of the two-dimensional pion multiplicity and invariant hadronic mass distributions.
Final-state interactions:
A new 16% uncertainty on the cross section for pion charge-exchange reactions for MeV was included [44]. As the new ND selections are sensitive to proton kinematics, an uncertainty was added that varies the fraction of events that include intranuclear nucleon rescattering while keeping the leptonic observables unmodified. The size of this uncertainty (30%) was motivated by a recent analysis of nuclear transparency data [45].
Near detector analysis— The neutrino flux and interaction models were constrained by fitting the unoscillated CC event spectra at the ND280. Three Time Projection Chambers (TPC) [46] with two Fine-Grained Detectors (FGD) [47] sandwiched between them track particles from interactions in the FGD target mass. Electromagnetic calorimeters surrounding the TPCs and FGDs, as well as the TPCs, are used to tag photons.
Selected CC events were separated into different samples according to the FGD in which the interaction occurred, the beam mode, the muon charge and the reconstructed final-state particle multiplicities. Negatively-charged muon candidates selected in -mode were divided into five samples per FGD: 1) at least one photon (CC-photon); 2) no photons, pions or protons (CC00p); 3) no photons or pions and at least one proton (CC0Np); 4) no photons and one positively-charged pion (CC1); 5) and all other CC events (CC-Other). Positively and negatively charged muon candidates in -mode are each divided into three samples per FGD: 1) no pions (CC0); 2) one pion with opposite charge to the muon (CC1); 3) and all other CC events (CC-Other). Positively charged muon candidates in -mode were not separately selected as the predicted contamination was 4%, compared to 30% negatively charged muon contamination in -mode [48].
Newly added proton-tagged samples offer enhanced abilities to constrain the -dependence of several uncertainties with data. Since ND280 achieves reliable proton tracking efficiency for momenta 450 MeV, samples with (without) protons generally correspond to higher (lower) average . The new photon-tagged sample helps to constrain production uncertainties, and increases the purity of other samples.
Each sample was binned according to the muon candidate’s momentum, , and the cosine of the angle between the muon momentum vector and the beam direction, . A MC prediction was made using the flux, cross-section and ND280 detector models, and an extended binned likelihood [49] fit to data was performed. Fig. 1 shows the data, prefit and postfit MC distributions for the FGD1 CC00p sample as a function of . The total -value is 0.575, indicating agreement between the data fit result and the input models. The uncertainty on the SK predicted event rates from cross-section and flux systematics was reduced from approximately 10–15% to 3–4% for each sample by using the ND280 constraint.
Figure 1: Data and model predictions before and after the ND280 fit for the -mode FGD1 CC00p sample binned in . The fit includes GeV for all ND280 samples, but the range is truncated here for readability. The postfit prediction is broken down by interaction channel. The predicted event rate from contamination in -mode is neglected in the figure as it only contributes at the sub-percent level.
SK event selection—Previous T2K analyses used five SK samples: 1 ring - and -like in both - and -modes and 1 ring -like with 1 decay electron in -mode [2] (tagging a pion below Cherenkov threshold). This analysis introduces a CC1-like sample in -mode, which tags pions through two topologies: (1) one ring each from a muon and a charged pion with one or two decay electrons (from the pion and muon decay); and (2) one -like ring (where the charged pion is below Cherenkov threshold) with two decay electrons (from the decay of the muon and the pion). Standard pre-selection criteria common to all SK samples [2] were applied in addition to the particle-identification (PID) requirements for the candidate -like, -like, and decay electron rings. These PID requirements reduce the number of selected background events with primary electrons or neutral pions, and those produced by neutral-current pion production processes.
The addition of the CC1-like sample increases the total number of selected -like events by 42.5%, although much of the increase is above the oscillation maximum, and is affected by different systematics to the dominant 1 ring -like samples. At a reference set of oscillation parameter values [50] (, , , eV2, eV2, and normal ordering), 53.5 (116.6) signal (total) events were predicted in this sample. A total of 134 data events passed the selection criteria with a total systematic uncertainty of 4.3%. The reconstructed neutrino energy for this sample, using only the reconstructed muon information, is
where , , are the proton, muon, and baryon rest masses; and , , are the muon-candidate reconstructed energy, three-momentum, and angle with respect to the neutrino beam. Predicted and observed distributions for this sample are presented in Fig. 2. Since events do not populate the region of the oscillation dip, we do not expect a significant increase in sensitivity to oscillation parameters. However, this sample provides improved control of background contributions and allows for valuable cross-checks of the cross-section model.
Figure 2: The distribution for the CC1 -mode SK sample shown for data and MC. The fit includes -like (-like) events with (1.25) GeV, but the range is truncated here for readability. The reference oscillation parameter values are described in the main text. The ND280 constraint has been applied to the MC prediction.
Oscillation Analysis— The PDG parameterization of the PMNS matrix [50] was used in the analysis. Matter effects were included with an average Earth crust density of 2.6 g/cm3. The peak neutrino energy and baseline used by T2K provide sensitivity to the PMNS parameters , , , and the magnitude and sign of the mass-splitting term .
Measurements were carried out with and without a Reactor Constraint (RC). The RC was included as a Gaussian prior on , the PDG world-average [51, 52, 53, 50]. Similarly, for and priors of and were used [50]. Flat priors were used for the other oscillation parameters. The impact of the choice of prior on was investigated and does not alter the conclusions of the analysis [54].
Two different statistical approaches were used to extract constraints on the oscillation parameters of interest from a likelihood function, the form of which remained unchanged from Ref. [2]. One used a Markov Chain Monte Carlo (MCMC) method [55, 56], to simultaneously fit the data from ND280 and SK, producing posterior distributions from which credible intervals were constructed. The other performed a piece-wise fit and marginalized over the parameters associated with the propagated ND280 constraint to provide frequentist confidence intervals, both with the constant and Feldman-Cousins (FC) [57] methods.
Tab. 1 presents the posterior predictive p-values [58, 59, 60] for all FD samples. After accounting for the ‘Look Elsewhere Effect’ using the Bonferroni correction [61], all samples pass the 5% threshold. This suggests that the model provides a plausible description of the data within the considered parameter space.
Sample
-value
1R-like
-mode
0.35
-mode
0.84
CC1-like
-mode
0.96
1R-like
-mode
0.13
-mode
0.63
1R-like 1d
-mode
0.89
Total
0.86
Table 1: Posterior predictive -values for every FD sample, from the fit including the RC.
Simulated Data Studies (SDS)— These study the impact of alternative interaction models and data-driven modifications of predictions at ND280 and SK, to check the completeness of our systematic uncertainty model, and are essential tests of the robustness of our result. Changes to and were deemed significant if the bias to the center of a 2 confidence interval was greater than 50% of the 1 Asimov interval width, or if the size of a one-dimensional, 2, FC-corrected interval changed by more than 10%. If a significant bias was found, additional smearing was applied to the relevant intervals. For , biases to interval boundaries which change the inclusion or exclusion of a physically interesting point in parameter space, e.g. , are reported as part of the result; but due to the non-Gaussian nature of the likelihood additional smearing was not applied to reported intervals.
SDS
HF CRPA
Bias
Size
Martini 1
Bias
Size
Non-CCQE
Bias
Size
SPP low-
Bias
Size
Table 2: Differences in the oscillation parameter constraints observed in the new and most impactful SDS. “Bias” shows changes to the center of the 2 confidence interval divided by the 1 Asimov interval width, and “size” is the change in size of the one-dimensional, 2, FC-corrected interval.
Various SDS were described in Ref. [2], and have also been carried out in this analysis. New SDS for this analysis are: replacing the default single-pion production model with the Martini et al. 1 model [62, 63] to test the robustness of the model for the CC1-like SK sample; altering the default nuclear response to the Hartree-Fock Continuum Random Phase Approximation (HF CRPA) [64, 65] to test the new proton-tagged ND280 samples. In total 19 SDS were performed. New SDS and those with the most significant impact are shown in Tab. 2. The observed bias on the center of the 2 interval exceeded the bias condition for both the HF CRPA and non-CCQE SDS. As a result, the contour was smeared by , determined from the quadrature sum of the biases on from all SDS. Additionally, a single SDS (SPP low- [40]) was found to change the confidence interval of such that is not excluded.
Figure 3: Credible regions in the – plane produced with the MCMC analysis, shown with and without the RC applied, and overlaid with the RC constraint from Ref. [50], for normal ordering.
Oscillation results and discussion—Fig. 3 shows credible regions in the – plane produced with and without the RC applied. The measurement of without the RC applied is consistent with the PDG value. When marginalized over both mass orderings, the best-fit value of with (without) the RC applied is ().
The FC-corrected frequentist confidence intervals for are shown in Fig. 4. The best-fit value is ( ) for normal (inverted) ordering with the RC applied. The data prefer values of close to radians, excluding values around radians at 3 in both orderings. The CP-conserving values of 0 and are excluded at the 3 level in inverted ordering. In normal ordering, is excluded at 90% confidence, and although the nominal fit excludes at 90% (Fig. 4), an SDS was found which could move the interval boundary past , so this value is not excluded in the reported result. The result was statistically limited and can be expected to improve as more data is accumulated.
Figure 4: The change in the best fit observed in the frequentist analysis as a function of and the mass ordering. Shaded regions correspond to the FC-corrected confidence intervals.Figure 5: Frequentist confidence intervals are shown in the – plane, produced using the constant method with the RC applied. The contour is smeared to cover the SDS as described in the main text.
Fig. 5 shows frequentist confidence intervals in the – plane for both mass orderings. The data exhibit a weak preference for the upper octant of and the normal ordering, with best-fit values of and . The MCMC analysis obtains a Bayes factor with (without) the RC applied of 2.3 (1.4) for the upper octant of over the lower and 2.7 (1.7) for the normal over inverted ordering.
The measurements presented thus far assume the PDG parameterization of the PMNS matrix. The Jarlskog invariant, , is a parameterization-independent way of measuring the scale of CP-violation generated by PMNS oscillations [66, 67]. A zero (non-zero) value for indicates CP-conservation (CP-violation) in three-flavor neutrino mixing. The constraint on , obtained with the MCMC analysis, and the impact of the choice of prior on that constraint, are shown in Fig. 6. The CP-conserving value, , is excluded at the 90% credible interval for both priors: flat in and flat in . Although changes in the prior were checked, the robustness of these credible intervals in has not been checked with SDS in this analysis.
Figure 6: Posterior probability distributions for the Jarlskog invariant taken from posterior distributions with priors that are either flat in (blue) or flat in (orange), obtained with the MCMC analysis for both orderings, with credible intervals (CIs) shown.
Conclusions— The T2K collaboration has measured the three-flavor PMNS neutrino oscillation parameters , , , , , and the mass ordering, using POT at SK. The analysis includes a new CC1-like SK sample; new ND event samples in -mode; significant improvements to the flux and neutrino interaction modeling; and a significantly expanded set of SDS to test the robustness of the analysis to out of model changes.
These improvements represent a significant step forward in the development of analysis methods which are robust to mis-modeling issues, a major obstacle for all current and future long-baseline neutrino oscillation experiments.
Our results show a weak preference for normal ordering and the upper octant of with best-fit values of and . One of the CP conserving values, , is excluded at 90% confidence, with a best-fit value of in normal ordering with the RC applied.
A data release including MCMC chains from the fit without the RC applied have been released in Ref. [68].
Acknowledgements—
The T2K collaboration would like to thank the J-PARC staff for superb accelerator performance. We thank the CERN NA61/SHINE Collaboration for providing valuable particle production data. We acknowledge the support of MEXT, JSPS KAKENHI and bilateral programs, Japan; NSERC, the NRC, and CFI, Canada; the CEA and CNRS/IN2P3, France; the Deutsche Forschungsgemeinschaft (DFG 397763730, 517206441), Germany; the NKFIH (NKFIH 137812 and TKP2021-NKTA-64), Hungary; the INFN, Italy; the Ministry of Science and Higher Education (2023/WK/04) and the National Science Centre (UMO-2018/30/E/ST2/00441, UMO-2022/46/E/ST2/00336 and UMO-2021/43/D/ST2/01504), Poland; the RSF (RSF 24-12-00271) and the Ministry of Science and Higher Education, Russia; MICINN (PID2022-136297NB-I00 /AEI/10.13039/501100011033/ FEDER, UE, PID2024-157541NB-I00 (UAM) and PID2023-146401NB-I00 (US), Severo Ochoa Centres of Excellence Programme 2025-2029 (CEX2024001442-S)), Government of Andalucia (FQM160) and the University of Tokyo ICRR’s Inter-University Research Program FY2025 Ref. J1, and ERDF and European Union (UAM: H2020-MSCA-RISE-GA872549- SK2HK) and NextGenerationEU funds (PRTR-C17.I1) and Generalitat de Catalunya (AGAUR 2021-SGR-01506, CERCA program), University of Seville grant (RYC2022-035203-I funded by MICIU/AEI/10.13039/501100011033, “ERDF a way of making Europe” and FSE+, Ayudas “Atracción de Investigadores con Alto Potencial” Ref. VIIPPIT-2025), and Secretariat for Universities and Research of the Ministry of Business and Knowledge of the Government of Catalonia and the European Social Fund (2022FI_B 00336), Spain; the SNSF and SERI, Switzerland; the STFC and UKRI, UK; the DOE, USA; and NAFOSTED (103.99-2023.144,IZVSZ2.203433), Vietnam. We also thank CERN for the UA1/NOMAD magnet, DESY for the HERA-B magnet mover system, the BC DRI Group, Prairie DRI Group, ACENET, SciNet, and CalculQuebec consortia in the Digital Research Alliance of Canada, and GridPP in the United Kingdom, the CNRS/IN2P3 Computing Center in France and NERSC, USA. In addition, the participation of individual researchers and institutions has been further supported by funds from the ERC (FP7), “la Caixa” Foundation, the European Union’s Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie grant, the Horizon Europe Marie Sklodowska-Curie Staff Exchange project JENNIFER3 grant 101183137; the JSPS, Japan; the Royal Society, UK; French ANR and Sorbonne Université Emergences programmes; the VAST-JSPS (No. QTJP01.02/20-22); and the DOE Early Career program, USA.
Abi et al. [2020]B. Abi et al. (DUNE), Deep Underground Neutrino Experiment (DUNE), Far Detector
Technical Design Report, Volume II: DUNE Physics, arXiv:2002.03005 [hep-ex] (2020).
Brun et al. [1994]R. Brun, F. Bruyant,
F. Carminati, S. Giani, M. Maire, A. McPherson, G. Patrick, and L. Urban, GEANT
Detector Description and Simulation Tool (1994).
Abgrall et al. [2016a]N. Abgrall et al. (NA61/SHINE), Measurements of ,
, , and proton production in
proton–carbon interactions at 31 GeV/c with the NA61/SHINE
spectrometer at the CERN SPS, Eur. Phys. J. C 76, 84 (2016a), arXiv:1510.02703 [hep-ex] .
Abgrall et al. [2016b]N. Abgrall et al. (NA61/SHINE), Measurements of
differential yields from the surface of the T2K replica target for incoming
31 GeV/c protons with the NA61/SHINE spectrometer at the CERN SPS, Eur. Phys. J. C 76, 617 (2016b), arXiv:1603.06774 [hep-ex] .
Abgrall et al. [2019]N. Abgrall et al. (NA61/SHINE), Measurements of , and proton double differential yields from the surface of the T2K replica
target for incoming 31 GeV/c protons with the NA61/SHINE spectrometer at the
CERN SPS, Eur. Phys. J. C 79, 100 (2019), arXiv:1808.04927 [hep-ex]
Ankowski et al. [2015]A. M. Ankowski, O. Benhar, and M. Sakuda, Improving the accuracy of neutrino
energy reconstruction in charged-current quasielastic scattering off nuclear
targets, Phys. Rev. D 91, 033005 (2015).
Chakrani et al. [2024]J. Chakrani et al., Parametrized uncertainties in the spectral function model of neutrino
charged-current quasielastic interactions for oscillation analyses, Phys. Rev. D 109, 072006 (2024), arXiv:2308.01838 [hep-ex] .
Rein and Sehgal [1981]D. Rein and L. M. Sehgal, Neutrino Excitation of
Baryon Resonances and Single Pion Production, Annals Phys. 133, 79 (1981).
Wilkinson et al. [2014]C. Wilkinson, P. Rodrigues, S. Cartwright, L. Thompson, and K. McFarland, Reanalysis of bubble
chamber measurements of muon-neutrino induced single pion production, Phys. Rev. D 90, 112017 (2014), arXiv:1411.4482 [hep-ex] .
Rodrigues et al. [2016]P. Rodrigues, C. Wilkinson, and K. McFarland, Constraining the GENIE
model of neutrino-induced single pion production using reanalyzed bubble
chamber data, Eur. Phys. J. C 76, 474 (2016), arXiv:1601.01888 [hep-ex]
Derrick et al. [1981]M. Derrick et al., Study of
single-pion production by weak neutral currents in low-energy
interactions, Phys. Rev. D 23, 569 (1981).
Radecky et al. [1982]G. M. Radecky et al., Study of
single-pion production by weak charged currents in low-energy
interactions, Phys. Rev. D 25, 1161 (1982).
Baker et al. [1981]N. J. Baker, A. M. Cnops,
P. L. Connolly, S. A. Kahn, M. J. Murtagh, R. B. Palmer, N. P. Samios, and M. Tanaka, Study of the isospin structure of single-pion production in
charged-current neutrino interactions, Phys. Rev. D 23, 2495 (1981).
Kitagaki et al. [1986]T. Kitagaki et al., Charged-current exclusive pion production in neutrino-deuterium
interactions, Phys. Rev. D 34, 2554 (1986).
Furuno et al. [2003]K. Furuno, A. Suzuki,
T. Kitagaki, M. Etou, H. Sagawa, K. B. McConnel Mahn, E. J. Jeon, and M. Sakuda, BNL
7-foot bubble chamber experiment: Neutrino deuterium interactions, in 2nd International Workshop on
Neutrino-Nucleus Interactions in the Few GeV Region (2003).
Aguilar-Arevalo et al. [2011a]A. A. Aguilar-Arevalo et al. (MiniBooNE), Measurement of Neutrino-Induced
Charged-Current Charged Pion Production Cross Sections on Mineral Oil at
E, Phys. Rev. D 83, 052007 (2011a), arXiv:1011.3572 [hep-ex] .
Aguilar-Arevalo et al. [2010]A. A. Aguilar-Arevalo et al. (MiniBooNE), Measurement of and
induced neutral current single production cross
sections on mineral oil at , Phys. Rev. D 81, 013005 (2010), arXiv:0911.2063 [hep-ex] .
Aguilar-Arevalo et al. [2011b]A. A. Aguilar-Arevalo et al. (MiniBooNE), Measurement of -induced
charged-current neutral pion production cross sections on mineral oil at
GeV, Phys. Rev. D 83, 052009 (2011b), arXiv:1010.3264 [hep-ex] .
Abe et al. [2020]K. Abe et al. (T2K), Measurement of the muon neutrino charged-current single
production on hydrocarbon using the T2K off-axis near detector
ND280, Phys. Rev. D 101, 012007 (2020), arXiv:1909.03936 [hep-ex] .
Niewczas and Sobczyk [2019]K. Niewczas and J. T. Sobczyk, Nuclear transparency in
monte carlo neutrino event generators, Phys. Rev. C 100, 015505 (2019).
Abe et al. [2017]K. Abe et al. (The T2K Collaboration), Measurement of neutrino and
antineutrino oscillations by the t2k experiment including a new additional
sample of interactions at the far detector, Phys. Rev. D 96, 092006 (2017).
Zyla et al. [2020]P. A. Zyla et al. (Particle Data Group), Review of Particle Physics, PTEP 2020, 083C01 (2020), and 2021
update.
Abe et al. [2014]Y. Abe et al. (Double Chooz), Improved measurements of the neutrino mixing
angle with the Double Chooz detector, JHEP 10, 086, [Erratum: JHEP 02, 074 (2015)], arXiv:1406.7763 [hep-ex] .
Bak et al. [2018]G. Bak et al. (RENO), Measurement of reactor antineutrino oscillation amplitude
and frequency at reno, Phys. Rev. Lett. 121, 201801 (2018), arXiv:1806.00248 [hep-ex] .
Abe et al. [2025a]K. Abe et al. (T2K), Testing T2K’s Bayesian constraints with priors in
alternate parameterisations, arXiv:2507.02101 [hep-ex] (2025a).
Metropolis et al. [1970]N. Metropolis, A. W. Rosenbluth, M. N. Rosenbluth, A. H. Teller, and E. Teller, Equation of state
calculations by fast computing machines, The Journal of Chemical Physics 21, 1087 (1970).
Gelman et al. [1996]A. Gelman, X.-L. Meng, and H. Stern, Posterior predictive assessment of model fitness
via realized discrepancies, Statistica sinica , 733 (1996).
Gelman et al. [2000]A. Gelman, Y. Goegebeur,
F. Tuerlinckx, and I. V. Mechelen, Diagnostic checks for discrete data
regression models using posterior predictive simulations, Appl. Statist. 49 Par 2, 247 (2000).
Martini et al. [2009] M. Martini, M. Ericson, G. Chanfray, and J. Marteau, A
Unified approach for nucleon knock-out, coherent and incoherent pion
production in neutrino interactions with nuclei, Phys. Rev. C 80, 065501 (2009), arXiv:0910.2622 [nucl-th] .
Jachowicz et al. [2002]N. Jachowicz, K. Heyde,
J. Ryckebusch, and S. Rombouts, Continuum random phase approximation approach to
charged current neutrino nucleus scattering, Phys. Rev. C 65, 025501 (2002).
Pandey et al. [2015]V. Pandey, N. Jachowicz,
T. Van Cuyck, J. Ryckebusch, and M. Martini, Low-energy excitations and quasielastic contribution to
electron-nucleus and neutrino-nucleus scattering in the continuum
random-phase approximation, Phys. Rev. C 92, 024606 (2015), arXiv:1412.4624 [nucl-th] .
Jarlskog [1985]C. Jarlskog, Commutator of the quark
mass matrices in the standard electroweak model and a measure of maximal
nonconservation, Phys. Rev. Lett. 55, 1039 (1985).
Krastev and Petcov [1988]P. I. Krastev and S. T. Petcov, Resonance Amplification
and t Violation Effects in Three Neutrino Oscillations in the Earth, Phys. Lett. B 205, 84 (1988).
Abe et al. [2025b]K. Abe et al. (The T2K Collaboration), Markov chain monte carlo chain release
for "results from the t2k experiment on neutrino mixing including a new far
detector muon-like sample", 10.5281/zenodo.17312780 (2025b).